You recieved an order to prepare 500ml of a 0.6% KCL solution. Your pharmacy only has a 10% KCL solution in stock. How many milliliters of the stock KCL solution would prepare this solution? How much distilled water water do you need to add to produce the final product?
Answer by stanbon(57332)

You received an order to prepare 500 ml of a 0.6% KCL solution. Your pharmacy only has a 10% KCL solution in stock. How many milliliters of the stock KCL solution would prepare this solution? How much distilled water do you need to add to produce the final product?
------------------------------------------------
Equation:
water + water = water
0.90x + x = 0.994*500
1.9x = 497
x = 261.58 (amt. of distilled water needed)
0.9x = 235.42 (amt of 0.6% KCL solution needed)
=================================================
Cheers,
Stan H.